Solution
= Solution
Let $A_n$ denote the event that at least one study is positive. <Conditional independence> gives $\mathbb P(A_n\mid T)=1-\beta^n$ and $\mathbb P(A_n\mid F)=1-(1-\alpha)^n$. Updating the <prior odds> by this event's <Bayes factor> gives
$$
\boxed{\frac{\mathbb P(T\mid A_n)}{\mathbb P(F\mid A_n)}
=R\frac{1-\beta^n}{1-(1-\alpha)^n}.}
$$
These are the <odds under at-least-one-positive selection>, not the odds using every study's actual result.